Grade 10Mechanics

200 ml of h202 having volume strength 22.7 is mixed with 800 ml h2o2 solution having volume strength 45.4 are mixed together. the density of final h2o2 solution is 1.0224 g/ml
find
1. final volume strength
2. molality of final solution
3. w/v % of final solution

SAMBHAV MISHRA
7 Years agoGrade 10

2 Answers

Nikhil
ApprovedApproved Tutor Answer7 Years ago
H2o2 ------  H20 + ½ O2 
volume strength of hydrogen peroxide is defined as amount of oxygen that can be produced by 1 litre of it...
Acc. New sign conventions 1mole gas = 22.7 ltr at STP
So Molarity of first sample is 2M and that of 2nd is 4M
so moles of h2o2 in 1st sample is 0.4 and in 2nd it is 4*800/1000 =3.2
after mixing voluume of final soln is1 ltr and total moles =3.6 this is also molarity 
so volume str = (3.6/2)*22.7 = 40.86 ltr (1.)
w/v% = weight of solute(h2o2)/volume of solution *100
         =3.6*34(molarmass of h2o2)/1000  *100
          =12.24% (3.)
molality = moles of solute/weight of solvent in kg
Now weight  of soln = density *volume = 1.0224*1000 =1022.4 gm
weight of h202 = 3.6*34= 122.4 gm
so weight of solvent = 1022.4-122.4 =900gm
             = 3.6/ 0.9 =4 (2.)
 
 
 
 
 
 
SAMBHAV MISHRA
7 Years ago
@nikhil thank you very much bro.. god bless you..................................u r doig a nice job 
keep it up...